# Mocap Everyone Everywhere: Lightweight Motion Capture With Smartwatches and a Head-Mounted Camera

### v1

@@ C.1 Sensors
Readings IMU Sensor Signals from Smartwatches. For demonstrations with smartwatches, we use Apple Watch SE2 [1] and the SensorLog app [8] to record and access IMU sensor data. From the recorded sensor signals, IMU rotations are obtained from “motionQuaternion” and accelerations from “motionAcceleration” of the Apple Watch OS [2]. The sensor data are recorded in 30 FPS. Furthermore, we apply an average filter on the acceleration data (window length: 7). As previously demonstrated [31], applying the average filter on both real and synthetic acceleration data makes the data sufficiently similar to each other. Note that our module is trained on synthetic IMU data, and we apply our trained model on real IMU data from smartwatches without additional fine-tuning, different from the previous methods [68, 67].
@@
Figure 8: (a) Placing IMU sensors for IMU sensor calibration. The negative z-axis of the sensors are aligned to the gravity direction, and the y-axis is set as a “facing direction”. (b) Placing cameras on axes $I_{x}$ and $I_{y}$ for camera coordinate alignment.
@@ C.2 Calibration and Alignment
IMU Sensor Calibration. In sensor data readings in real IMU sensors (e.g., smartwatches), calibration should be done beforehand as the sensor data are in arbitrary coordinates. The IMU sensor data should first be calibrated to be coherent with the real-world coordinates where the gravity direction is in the negative z-axis. Here the y-axis is defined as the user’s facing direction. We use the notations similar to TIP [31], which are as follows:
@@
Let the arbitrary coordinate where raw IMU sensors are defined as $r$, and the global coordinate coherent with real-world coordinates as $g$. Finally, as the module $\mathcal{F}_{est}$ expects wrist joint rotations, the rotation value of IMU sensors should be calibrated into wrist joint rotations $\mathbf{R}_{g}^{j}$. Similarly, the raw acceleration data are defined in a sensor-local coordinate and should be converted into coordinate $g$.
$\mathbf{R}_{r}^{S}$ indicate rotations of raw IMU sensor data defined in the initial arbitrary coordinate. At the first calibration stage, the user is asked to align the smartwatches to the specified coordinate $g$ (Fig. 8 (a)) so that in the moment $\mathbf{R}_{r}^{S}=\mathbf{R}_{r}^{g}$. $\mathbf{R}_{g}^{r}$ is derived from $(\mathbf{R}_{r}^{g})^{-1}$. Following previous methods [31, 68], the sensors are placed still for 3 seconds and the signal is averaged. Provided that the axes are aligned, it is not necessary for the two IMU sensors to be in the exact same position as IMU sensor signals are translation invariant.
@@
Synchronization. The initial time synchronization between IMU sensors and the head-mounted camera is set by the timecodes generated from the sensors. However, empirically we found there were minor errors in the timecode and therefore should be adjusted. For the adjustment, the user is asked to stand still and swing their arms up and down, as in Fig. 10 (a). When the hands move up to the highest position close to the camera (Fig. 10 (a)), the IMU sensor signals show a specific peak as in Fig. 10 (b). The time synchronization is done by aligning the two. We empirically found out that swinging is more robust than clapping, as in some cases the force transmitted to the wrist sensor during clapping can cause sensor peaks to spike.
@@ D.4 Motion Optimization
Models and Training. The encoder and decoder of the autoencoder structure to build motion manifolds (Sec 3.4) consist of 3 layers of 1D temporal-convolutions with a kernel width of 25 and stride 2. The channel dimension of each output feature is set to 256. The autoencoder is trained with the AdamW optimizer [40] and the learning rate is automatically set with DAdaptation [16].For optimizing the latent vector within the manifold, Adam optimizer [35] was used, and the learning rate was set to 0.0007 for single-person egocentric visual cues $\phi_{E}$, 0.001 for multi-person visual cues $\phi_{T}$. As in the motion estimation module, we use PyTorch [50] and FairMotion library [22] for implementation. Nvidia RTX 3090Ti GPU was used for training and optimization. The datasets used for training the autoencoder are identical to the datasets used in training $\mathcal{F}_{est}$ (Supp. Sec. D.3).

### v2

@@ C.1 Sensors
Readings IMU Sensor Signals from Smartwatches. For demonstrations with smartwatches, we use Apple Watch SE2 [1] and the SensorLog app [8] to record and access IMU sensor data. From the recorded sensor signals, IMU rotations are obtained from “motionQuaternion" and accelerations from “motionAcceleration" of the Apple Watch OS [2]. The sensor data are recorded in 30 FPS. Furthermore, we apply an average filter on the acceleration data (window length: 7). As previously demonstrated [31], applying the average filter on both real and synthetic acceleration data makes the data sufficiently similar to each other. Note that our module is trained on synthetic IMU data, and we apply our trained model on real IMU data from smartwatches without additional fine-tuning, different from the previous methods [68, 67].
@@
data rep multi-stage MPJPE($\downarrow$) r.MPJPE($\downarrow$) MPJVE($\downarrow$)
8.50 5.85 27.63
$\checkmark$ 6.71 5.58 23.69
$\checkmark$ $\checkmark$ 5.20 4.95 17.00
Table 7: Ablation studies on the components of $\mathcal{F}_{est}$. Settings are identical as in Tab. 2 (AMASS) in the main paper.
@@ C.2 Calibration and Alignment
IMU Sensor Calibration. In sensor data readings in real IMU sensors (e.g., smartwatches), the IMU sensor data should first be calibrated to be coherent with the real-world coordinates where the gravity direction is in the negative z-axis. Here the y-axis is defined as the user’s facing direction. We use the notations similar to TIP [31], which are as follows:
@@
Let the reference coordinate where raw IMU sensors signals are defined as $r$, and the global coordinate coherent with real-world coordinates as $g$. Finally, as the module $\mathcal{F}_{est}$ expects wrist joint rotations, the rotation value of IMU sensors should be calibrated into wrist joint rotations $\mathbf{R}_{g}^{j}$. Similarly, the raw acceleration data are defined in the reference coordinate of sensors and should be converted into coordinate $g$. Fig. 10 includes visualization of raw and calibrated signals.
$\mathbf{R}_{r}^{S}$ indicate rotations of raw IMU sensor data defined in the initial reference frame. At the first calibration stage, the user is asked to align the smartwatches to the specified coordinate $g$ (Fig. 8 (a)) so that in the moment $\mathbf{R}_{r}^{S}=\mathbf{R}_{r}^{g}$. From $\mathbf{R}_{r}^{g}$ we can derive $\mathbf{R}_{g}^{r}$ by applying $(\mathbf{R}_{r}^{g})^{-1}$. Following previous methods [31, 68], the sensors are placed still for 3 seconds and the signal is averaged. Provided that the axes are aligned, it is not necessary for the two IMU sensors to be in the exact same position as IMU sensor signals are translation invariant.
@@
Figure 8: (a) Placing IMU sensors for IMU sensor calibration. The negative z-axis of the sensors are aligned to the gravity direction, and the y-axis is set as a “facing direction". (b) Placing cameras on axes $I_{x}$ and $I_{y}$ for camera coordinate alignment.
@@
Figure 10: Visualization of the raw and calibrated IMU signals defined in reference coordinate $r$ and user-specified global coordinate $g$.
@@
Synchronization. The initial time synchronization between IMU sensors and the head-mounted camera is set by the timecodes generated from the sensors. (For the camera, the timecode is recorded when the shutter is activated) However, empirically we found there were minor errors in the timecode and therefore should be adjusted. For the adjustment, the user is asked to stand still and swing their arms up and down, as in Fig. 11 (a). When the hands move up to the highest position close to the camera (Fig. 11 (a)), the IMU sensor signals show a specific peak as in Fig. 11 (b). The time synchronization is done by aligning the two. To adjust between the camera and the IMU sensors, the frame where the hands are closest to the camera is selected. We empirically found out that swinging is more robust than clapping, as in some cases the force transmitted to the wrist sensor during clapping can cause sensor peaks to spike.
C.3 Capture Setup for Ablations
For quantitative evaluation of ablative baselines (Sec. 4.4) XSens MVN Link [10] was used to capture ground-truth data. As an IMU-based method, however, XSens also suffers from root drift issues [23]. We ignore these factors during the evaluation by considering errors in XY aligned space (in evaluating floor update algorithm) or capturing motion with relatively small root translations (in evaluating $\mathcal{F}_{opt}$).
For the ablation of floor plane update, the height changes from 0m to --4.21m. For ablations on visual-cue based motion optimization $\mathcal{F}_{opt}$ in egocentric cases, we capture 3 scenarios (making coffee, using coffee machine, taking snacks off the shelf; total 2289 frames) and the right hand was detected as the visual cue $\phi_{E}$.
@@ D.4 Motion Optimization
Models and Training. The encoder and decoder of the autoencoder structure to build motion manifolds (Sec 3.4) which consist of 3 layers of 1D temporal-convolutions with a kernel width of 25 and stride 2. The channel dimension of each output feature is set to 256. The autoencoder is trained with the AdamW optimizer [40] and the learning rate is automatically set with DAdaptation [16].For optimizing the latent vector within the manifold, Adam optimizer [35] was used, and the learning rate was set to 0.0007 for single-person egocentric visual cues $\phi_{E}$, 0.001 for multi-person visual cues $\phi_{T}$. As in the motion estimation module, we use PyTorch [50] and FairMotion library [22] for implementation. Nvidia RTX 3090Ti GPU was used for training and optimization. The datasets used for training the autoencoder are identical to the datasets used in training $\mathcal{F}_{est}$ (Supp. Sec. D.3).